OpenStudy (aaronandyson):

A moving body of mass 400N posses 500J of Kinetic Energy.Calculate it's velocity take g = 10ms^-2 @perl @rvc

rvc (rvc):

What are things are given? Make a note of the data

OpenStudy (aaronandyson):

Mass=400N K.E=500J g = 10ms^-2 v = ?

rvc (rvc):

correct :) Now what is the formula of KE ?

OpenStudy (aaronandyson):

K.E = mv^2/2

rvc (rvc):

correct :) we need velocity so keep the v term on the RHS and the rest on the LHS

OpenStudy (aaronandyson):

But,Mass gievn 400N I need to convert it to kg

rvc (rvc):

yeah do you know the conversion?

OpenStudy (aaronandyson):

No :(

rvc (rvc):

oh well what is the relation between newton and kg?

OpenStudy (aaronandyson):

iDK Newtonis unit of force while kg is unit of mass :/

rvc (rvc):

mg

OpenStudy (anonymous):

\[ma=mv ^{2}/2\]

rvc (rvc):

yeah here for a we put g

OpenStudy (anonymous):

mass gets cut off... hence dont worry abt mass

rvc (rvc):

nope wait we have KE=500J we need to convert mass

OpenStudy (anonymous):

oh ya... sry.. i messed up..

OpenStudy (aaronandyson):

I'm confused :(

rvc (rvc):

well u just need to convert N to kg and rest is easy divide the mass by 10

OpenStudy (aaronandyson):

1N = how many kg? :/

OpenStudy (aaronandyson):

@rvc

rvc (rvc):

sorry os is lagging for me

rvc (rvc):

1N=0.10kg

rvc (rvc):

just divide the mass by 10

OpenStudy (aaronandyson):

How? How 10?

rvc (rvc):

oops

rvc (rvc):

multiply

rvc (rvc):

@IrishBoy123 please confirm

OpenStudy (aaronandyson):

im confused :/

rvc (rvc):

i just want to confirm about the conversion @Callisto

OpenStudy (aaronandyson):

HELP :(

OpenStudy (aaronandyson):

@butterflydreamer

OpenStudy (aaronandyson):

@mathmate

rvc (rvc):

@Michele_Laino will surely help :)

OpenStudy (michele_laino):

hint: if you use the Newton as unit of measure, then 400N is a weight, not a mss

OpenStudy (michele_laino):

oop.. ...not a mass...

rvc (rvc):

so its F=mg therefore m= \(\frac{f}{g}\)

OpenStudy (michele_laino):

that's right! @rvc So, m=400/9.81=...? @AaronAndyson

OpenStudy (michele_laino):

or: m=400/10=...?

rvc (rvc):

well it is given that g=10m/s^2

rvc (rvc):

haha so you need to divide by 10

OpenStudy (aaronandyson):

40? But,I still can't understand HOW?

rvc (rvc):

what is the formula for force???

OpenStudy (aaronandyson):

F = MG

OpenStudy (michele_laino):

ok! our particle has a mass of 40 Kg so its sped is given by the subsequent formula: \[\Large v = \sqrt {\frac{{2KE}}{m}} = ...?\] where KE is its kinetic energy

rvc (rvc):

correct

rvc (rvc):

I think he was confused with the conversion part only well rest is easy just put those values and find out the answer

OpenStudy (michele_laino):

\[\Large v = \sqrt {\frac{{2 \times KE}}{m}} = \sqrt {\frac{{2 \times 500}}{{40}}} = ...?\]

OpenStudy (michele_laino):

what is v? @AaronAndyson

rvc (rvc):

All the best!! @AaronAndyson

OpenStudy (aaronandyson):

5?

OpenStudy (michele_laino):

that's right! v= 5 meters/second

OpenStudy (aaronandyson):

Yet,I have no idea HOW

OpenStudy (michele_laino):

keep in mind that the kinetic energy KE, of a particle, whose mass is m, and whose speed is v, is given by the subsequent formula: \[\Large KE = \frac{1}{2} \times m \times {v^2}\]

OpenStudy (michele_laino):

since the weight of your particle, is P=400 Newton, then its mass is: \[\Large m = \frac{P}{g} \cong \frac{{400}}{{10}} = 40\;{\text{Kg}}\]

OpenStudy (aaronandyson):

DUH Thanks LOL XD

OpenStudy (michele_laino):

thanks!